Please show all work to fully understand a. Given that there is a recessionary gap of 400 Billion dollars and we have an MPC = 0.75, calculate the amount of spending that would be necessary to achieve FE-GDP? Hint: ΔGDP- 400B and you want to find the change in AE. Use the equations below. (0) Exp. Multiplier 1-MPC (ii)Δ GDP : Exp Mult.AAE b. Using the calculations above show the results on the AE and AD/AS graphs. Make sure to use your calculations when labeling the graphs. c. How would your answer change if instead the Federal Government wanted to achieve the same result through taxes? Hint: This is a calc-) Δtax-MAE MPC
Answer #1

a) If MPC = 0.75, expenditure multiplier = 1/1-MPC = 1/1-0.75 = 4. GDP gap is 400 billion and government want to eliminate it. Hence the required change in AE is

change in Y/Change in AE = Multiplier

400 B/ change in AE = 4

Change in AE = 400/4 = 100 billion

Government spending should be increased by 100 billion

b) Graph is shown below

Actual expenditure Expenditure Planned Δ G-100B expenditure ΔΥ-400B 45。 Y2 Income, output, Y

AS AD Real output

c) Tax multiplier = -MPC/1-MPC = -0.75/0.25 = -3

GDP gap is 400 billion and government want to eliminate it by reducing taxes. Hence the required change in AE is

change in Y/Change in AE = tax multiplier

400 B/ change in AE = -3

Change in AE = 400/-3 = -133.33 billion

Hence taxes should be reduced (due to negative sign) by 133.33 billion.